Depot report - October 10th/11th 2020

Mike Hallam-Rudd writes:

Another weekend gone and more ticked off at the DPS Depot. Bob continued with the PU Feet with help from Graham in finding the correct style 'lock nuts' for the main bolt which goes through the foot and onto the foot of the PU itself. Apart from the topcoat which will be done when they are fitted to 15 this task is now complete. Howard, with daughter Mia, set about clean down the cubicle doors for 15's No 1 cab along with the back panels for the instrument dials etc. These have all been painted previously but have attracted dirt whilst in storage, now ready for fitting. Mike, with assistance from Howard and supervised by Geoff (Duty RO with bad back) fitted the roof covers over 19's cooler groups as it will be outside for a further 3 weeks whilst a reshuffle of the stores area adjacent to No 1 road where it will be parked over the winter.....and ready to go as soon as we get the all clear. Mike then cleared a few more items out of the way ready for the return of No 8's cab on Wednesday 14th courtesy of Merritts Machine Movers and their 24 ton fork lift after a project carried out by Crowle Wharf Engineering, Crowle, Nr Scunthorpe. Look out for it on the M18 jtn 5 to 1 then M1 jtn 32 to 30 mid-morning on the day.

Both Steves were working on 15's bodywork. Steve P continued with his ongoing work inside whilst Steve Mac attended to some adjustments to the holes for fixing a bodyside window on B side No 2 end. Pete and Dave did some more work on PU 451. The engineering team have decided to try and recover the generator fitted to power unit 417 so this was also done by Pete and Dave. This is an effort to try to get a true mega reading on the generator as it has stood in the depot for nearly 10 years now may have attracted some dampness. The generator is now covered with old but unused pieces of PU lagging to retain the heat. This process is to identify if this generator has the right readings, if it does we may be able to revarnish the armature ourselves thus saving many thousands of pounds which we honestly do not have at the moment.

Graham finalised his calculations working out how much cable and fittings he needs to purchase for the replacement power cables on D9009. Alex continued his painting project on No 9's bogie and buffer beams. Before we all left one of the 2 new heaters was turned 90 deg and then carefully lifted from its 2 joined up pallets and put in its final resting place. We are just awaiting for the local flue contractor to come on site and give us a quote for the supply and installation of the flue.....if anybody out there is in the industry and wishes to have a look at the job please contact me at and I will pass your details to Dave who is leading this project to keep the volunteers and locos warm through the winter.

Photos attached copyright Alex Stojanovic